# Evaluating multiple models using labeled and unlabeled data

### v2

@@ 5.1 Datasets and Classifier Sets
We select datasets and classifier sets to be realistic and diverse, capturing multiple modalities (EHRs, text, graphs, and images), domains (healthcare, content moderation, chemistry), and architectures (spanning logistic regressions to large language models). We report ground truth metrics for the binary and multiclass classifiers in Tables S1 and S2 respectively, and a detailed description of each dataset and classifier set in Appendix B.1. We summarize each dataset and differences between classifiers in the associated classifier set below.
We test on six different classification datasets: (1) MIMIC-IV (Johnson et al., 2020), a dataset of Boston-area medical records with three hospitalization outcomes (critical outcome, ED revisit within 30 days, and hospital admission) with classifiers drawn from prior works (Movva et al., 2023), (2) CivilComments (Borkan et al., 2019), a dataset of social media comments flagged as “toxic” or not by human annotators, with pretrained classifiers downloaded from the WILDS benchmark (Koh et al., 2021), (3) OGB-SARS-CoV (Hu et al., 2020), a molecular property prediction task, with pretrained classifiers downloaded from WILDS, (4) MultiNLI, a a natural language task, with classifiers drawn from the SubpopBench benchmark (Yang et al., 2023), (5) ImageNetBG, a modified version of ImageNet with nine coarse-grained classes, with classifiers drawn from SubpopBench, and (6) AG News (Zhang et al., 2015), a news article classification task, using fine-tuned open-sourced LLMs available on HuggingFace as our classifiers.
@@ 5.2 Baselines
We compare against seven baselines that (with the exception of Labeled) make use of both labeled and unlabeled data to arrive at performance estimates:
  1. 1.
    Labeled represents the standard approach to classifiers evaluation and compares classifier scores to ground truth labels, only using examples for which labels are available.
  2. 2.
    Pseudo-Labeled trains a logistic regression classifier to predict the true label from the classifier scores—directly estimating $P(y^{(i)}|\mathbf{s}^{(i)})$—and then labels the unlabeled examples using this classifier.
  3. 3.
    Majority-Vote ensembles the classifier predictions by performing an accuracy-weighted aggregation of each classifier’s prediction. We assign weights to each classifier based on the accuracy achieved on the small sample of labeled data.
  4. 4.
    Dawid-Skene (Dawid & Skene, 1979), uses multiple noisy discrete annotations to estimate the latent true label of each example. Dawid-Skene assumes class-conditional errors are independent across annotators (i.e. $P(\hat{y}\neq y|y=k)$ is independent).
  5. 5.
    Bayesian Calibration (Ji et al., 2020) re-calibrates a classifier’s predicted probabilities based on the observation that if a classifier’s scores are calibrated, estimating performance from the predictions alone is possible. Bayesian Calibration does not extend to multiclass settings, so we can only use this baseline in our binary experiments.
  6. 6.
    AutoEval (Boyeau et al., 2024) learns a rectifier on limited labeled data to debias classifier predictions. This rectifier is then applied to the unlabeled examples to estimate metrics of interest. We compare to AutoEval on accuracy; additional metrics are not directly supported by the public implementation.
  7. 7.
    Active-Testing (Kossen et al., 2021) reframes semi-supervised evaluation as an active learning problem, in which a method selects which examples to label out of a large pool of unlabeled examples. The performance metric is then computed according to the sample of labeled data. We compare to Active-Testing on accuracy, a metric for which an acquisition strategy is available.
For baseline implementation details, please refer to Appendix B.2. We also provide a comparison to ensembling (i.e., drawing the “ground truth” label based on the average classifier score) in Appendix D.4, and a discussion of connections to weak supervision in Appendix D.3.
@@ 5.3 Evaluation
We evaluate SSME’s ability to estimate four continuous performance metrics for each binary classifier: accuracy, area under the receiver operating characteristic curve (AUC), area under the precision-recall curve (AUPRC), and the expected calibration error (ECE). For multi-class problems, we evaluate accuracy and top-label calibration error (Gupta & Ramdas, 2022); evaluations of multiclass AUC and AUPRC are also possible, but less standard.
We partition each dataset into three splits: the classifier training split (which we use to train the classifiers whose performance we will estimate), the estimation split (which we use to fit SSME and estimate classifier performance), and the evaluation split (which we use for a held-out, ground-truth measure of classifier performance). All splits are sampled from the same distribution, except when estimating subgroup-specific performance, where the evaluation split pertains to a single subset of the test distribution.
To evaluate metric estimates, we measure the absolute error of the estimated metric, computed using the estimation split, compared to the true metric, computed on the held-out evaluation split (averaging over classifiers in the set). The estimation split consists of either 20, 50, or 100 labeled examples and 1000 unlabeled examples across all experiments. The size of the evaluation split is on the order of thousands of labeled examples and varies by task (see Appendix B.1 for exact split sizes). For each task, we report results over 50 random samples of the splits. In line with prior work, we report rescaled estimation error for each metric (where all errors are relative to using labeled data alone), allowing us to standardize the scale of errors across datasets and metrics (Garg et al., 2022).
@@ 6 Results
6.1 Classifier evaluation on binary tasks
We now compare SSME to seven baselines in terms of its ability to estimate classifier performance on five binary tasks. All figures report rescaled metric estimation error (RMAE; lower is better) and reflect performance estimation using 20 labeled examples and 1000 unlabeled examples. Rescaling metric estimation error allows us to aggregate performance across tasks and metrics, and is standard in prior work (Garg et al., 2022). Results are consistent across additional values of $n_{\ell}$ (50 and 100; see results in D.1), although labeled data grows more competitive (as expected) with larger labeled dataset sizes.
Figure 2: Metric estimation error on binary tasks ($\mathbf{n_{\ell}=20}$, $\mathbf{n_{u}=1000}$). Each point plots the rescaled mean absolute error (RMAE) across runs, where 1.0 (dashed line) is the RMAE of using labeled data alone. SSME (gray) achieves lower estimation error than do the baselines (averaging across metrics and 50 runs) and reported across 5 binary tasks (y-axis). Results reporting mean absolute error (unscaled) and standard deviations are in Tables 1, S3, S4, and S5.
Dataset $n_{\ell}$ $n_{u}$ Labeled Majority-Vote Pseudo-Labeled Dawid-Skene AutoEval Active-Testing Bayesian-Calibration SSME-KDE-M SSME-KDE (Ours)
Critical Outcome 20 1000 5.19 ± 3.85 4.92 ± 0.21 4.12 ± 3.87 4.36 ± 0.31 4.78 ± 3.34 5.17 ± 1.16 2.80 ± 2.23 1.70 ± 0.99 0.67 ± 0.46
50 1000 2.90 ± 2.13 4.71 ± 0.23 3.06 ± 2.32 4.07 ± 0.40 3.01 ± 2.36 5.61 ± 1.44 2.07 ± 1.29 1.65 ± 0.90 0.78 ± 0.47
100 1000 2.09 ± 1.47 4.55 ± 0.31 1.58 ± 1.08 3.87 ± 0.38 2.00 ± 1.16 5.48 ± 1.25 1.18 ± 0.74 1.30 ± 0.70 0.77 ± 0.47
ED Revisit 20 1000 5.11 ± 3.53 2.15 ± 0.08 5.13 ± 3.23 4.02 ± 2.83 4.70 ± 3.32 2.83 ± 0.71 4.36 ± 2.76 1.64 ± 1.24 0.45 ± 0.36
50 1000 2.02 ± 2.08 2.10 ± 0.11 2.73 ± 2.24 2.74 ± 2.22 1.95 ± 2.07 3.03 ± 0.96 2.47 ± 2.07 1.46 ± 0.97 0.53 ± 0.39
100 1000 1.43 ± 1.15 2.01 ± 0.15 1.54 ± 1.22 1.51 ± 1.18 1.42 ± 1.04 2.64 ± 0.57 1.43 ± 1.12 1.18 ± 0.89 0.57 ± 0.39
Hospital Admission 20 1000 7.32 ± 4.52 19.68 ± 0.44 6.86 ± 4.31 19.55 ± 0.47 7.19 ± 3.73 9.33 ± 3.03 2.48 ± 1.59 3.29 ± 1.71 1.88 ± 1.04
50 1000 5.40 ± 2.98 18.91 ± 0.51 3.99 ± 2.97 18.78 ± 0.51 5.23 ± 2.46 9.25 ± 2.32 2.14 ± 1.28 3.17 ± 1.85 1.95 ± 0.99
100 1000 3.64 ± 1.99 18.02 ± 0.52 3.01 ± 1.92 17.81 ± 0.59 4.01 ± 1.99 9.24 ± 2.90 2.42 ± 1.19 3.06 ± 1.64 1.51 ± 0.82
SARS-CoV Inhibition 20 1000 6.11 ± 3.45 5.44 ± 0.21 5.95 ± 3.62 4.91 ± 0.65 4.59 ± 3.05 5.97 ± 1.69 2.25 ± 1.13 3.06 ± 0.83 2.30 ± 0.56
50 1000 3.22 ± 2.05 5.33 ± 0.23 2.99 ± 1.64 4.50 ± 0.63 2.64 ± 1.53 5.74 ± 1.14 1.74 ± 0.76 2.59 ± 0.94 2.35 ± 0.35
100 1000 2.04 ± 1.38 5.07 ± 0.27 2.14 ± 1.10 4.01 ± 0.62 1.94 ± 0.90 5.99 ± 1.70 1.43 ± 0.68 1.84 ± 0.85 2.36 ± 0.47
Toxicity Detection 20 1000 5.95 ± 2.64 4.62 ± 0.31 5.03 ± 2.91 4.82 ± 0.32 5.27 ± 2.71 7.19 ± 1.57 5.29 ± 1.06 6.71 ± 0.83 2.34 ± 0.52
50 1000 4.03 ± 2.44 4.47 ± 0.29 2.88 ± 1.72 4.65 ± 0.29 3.37 ± 1.48 7.26 ± 1.71 4.57 ± 1.07 5.38 ± 1.01 2.22 ± 0.47
100 1000 2.43 ± 1.48 4.26 ± 0.40 1.90 ± 1.11 4.46 ± 0.40 2.34 ± 0.94 7.50 ± 2.16 3.78 ± 0.92 3.80 ± 1.16 2.14 ± 0.54
Table 1: Mean absolute error in accuracy estimation on binary tasks. We report mean absolute error (averaging across classifiers) across five binary classification tasks and different amounts of labeled data. We bold the best performing method in each row, and underline the next best performing method. SSME-KDE-M describes performance when fitting SSME to a single classifier’s scores, instead of modeling the joint distribution of $p(y,s)$.
@@ 6.1 Classifier evaluation on binary tasks
SSME achieves lower mean estimation error (averaging across tasks and metrics) than all baselines, indicating more accurate estimation of classifier performance. Concretely, SSME reduces estimation error by 5.1$\times$ relative to labeled data alone (averaged across tasks and metrics). In contrast, the next best method reduces estimation error by 2.4$\times$. SSME also outperforms baselines on specific metrics. For accuracy, SSME reduces metric estimation error, relative to using labeled data alone, by 5.6$\times$ (averaged across tasks); the next best method for each dataset reduces metric estimation error by 2.0$\times$. While the magnitude by which SSME beats baselines varies —for example, SSME reduces error by 2.9$\times$ on AUC (averaged across tasks), while the next best method reduces error by 2.6$\times$ — SSME consistently outperforms baselines across metrics.
Our results are also encouraging in absolute terms. With 20 labeled examples and 1000 unlabeled examples, SSME estimates accuracy within 1.5 percentage points (averaging across tasks). The closest baseline estimates accuracy within 3.4 percentage points. Results comparing SSME to the next best baseline on other metrics (1.9 vs 3.8 on ECE; 3.6 vs 4.3 on AUC; 8.5 vs 10.2 on AUPRC) confirm that SSME not only achieves more accurate classifier performance estimation compared to prior work, but that the resulting performance estimates are reasonably close to the ground truth measurements for each metric.
SSME estimates performance more accurately because it makes use of all available information: multiple classifiers, continuous scores over all classes, and unlabeled data. Dawid-Skene [^1] Fig. 2 omits Dawid-Skene from accuracy results on hospital admission because its much higher RMAE of 2.6 distorts the plotting scale. and AutoEval discretize classifier scores (although AutoEval does make use of classifier confidence associated with each discrete prediction). Pseudo-Labeling, Bayesian-Calibration, and AutoEval each learn a mapping from s to $y$ using only the labeled data and apply that mapping to the unlabeled data (rather than learning from labeled and unlabeled data together). By jointly learning across both labeled and unlabeled data, SSME is able to generalize much better in cases where there aren’t enough labels to estimate the joint distribution of classifier scores and labels from labeled examples alone. Finally, Bayesian-Calibration learns from a single classifier’s scores, and does not learn from multiple models at once.
@@
SSME provides the greatest benefits relative to labeled data alone when measuring expected calibration error (ECE), with a reduction in estimation error of 7.2$\times$ (averaging across tasks). ECE is harder to estimate with few labeled examples because it requires binning and then averaging calibration error across bins. This process tends to yield greater variability when the number of labeled points per bin is small. We observe the smallest benefits relative to labeled data alone when measuring AUPRC (a reduction in estimation error of 2.2$\times$, relative to labeled data).
@@
SSME’s performance continues to improve with more labeled data, but the advantage it confers over labeled data decreases: for example, with 20, 50, and 100 datapoints, SSME outperforms labeled data alone by 5.6$\times$, 3.0$\times$, and 1.6$\times$. Similar to labeled data alone, there are diminishing but positive returns to adding labeled data to SSME’s performance estimation procedure.
Another way to quantify SSME’s benefit is to measure the amount of labeled data required to match SSME’s performance, or the effective sample size (ESS), as introduced by prior work (Boyeau et al., 2024) (see Appendix B.4 for implementation details). With access to 20 labeled exampels and 1000 unlabeled examples, SSME achieves an average ESS of 539 labeled examples for estimating ECE (averaging over tasks). In contrast, the next best approach achieves an ESS of 110 labeled examples to estimate ECE.
@@
To validate the benefit of fitting the mixture model to multiple classifiers simultaneously, we compare to an ablated version of SSME fit on a single classifier at a time. In this setting, SSME estimates the classifier-specific marginal distribution of $P(y|\textbf{s})$, and uses this estimate to evaluate the classifier in question. Doing so results in worse performance estimates across metrics, tasks, and amounts of labeled data (Tables 1, S3, S4, S5) relative to our full model. This agrees with findings from the ensembling literature (Schapire, 1990): each classifier provides distinct information about the ground truth label for a given example, which SSME is able to use.
6.2 Classifier evaluation on multi-class tasks
Figure 3: Metric estimation error on multiclass tasks ($\mathbf{n_{\ell}=20}$, $\mathbf{n_{u}=1000}$). SSME-NF consistently reduces estimation error relative to labeled data alone, and more consistently than any baseline that is able to estimate performance in multiclass settings.
We now validate SSME’s performance on multi-class tasks; Figure 3 reports our results. Because the utility of kernel density estimators is known to degrade in higher dimensional problems (as can be the case with multi-class outputs)  (Jiang, 2017), we provide an instantiation of SSME using a normalizing flow, which we term SSME-NF (for additional implementation details, see Appendix C). Normalizing flows have been shown to effectively model mixtures of high-dimensional distributions (Izmailov et al., 2020); our results align with these findings. SSME-NF improves over labeled data by 2.9$\times$ (averaging over tasks and metrics); SSME-KDE improves over labeled data by 2.7$\times$. In contrast, the next best baseline (Dawid-Skene) improves over labeled data by 1.5$\times$. On each dataset and metric, SSME-NF is consistently one of the top two methods. Our results are encouraging in the context of evaluating large language models as annotators, an emerging application (Ziems et al., 2024). The classifier sets we evaluate on MultiNLI and AG News reflect real multi-class settings where a practitioner might wish to evaluate several language model based classifiers: the MultiNLI classifier set contains off-the-shelf language model based classifiers, while the AG News classifier set contains the top 10 most-downloaded sentence encoders on HuggingFace, fine-tuned on 200 labeled examples using SetFit (Tunstall et al., 2022).
6.3 Assessing subgroup-specific performance
@@
Bayesian-Calibration 0.45 0.33 0.41
SSME-KDE (Ours) 0.42 0.19 0.39
@@
Table 2: Subgroup-specific performance estimation ($\mathbf{n_{l}=20}$, $\mathbf{n_{u}=1000}$). SSME achieves the lowest rescaled metric estimation error (RMAE) (averaging across metrics and demographic subgroups).
SSME can be applied to measure performance within demographic groups of interest, a task central to assessments of algorithmic fairness (Chen et al., 2021). These groups may be based on gender, age, or race, or other demographic variables. Unlabeled data has been used in multiple ways to improve subgroup evaluation and algorithmic fairness (Ji et al., 2020; Sagawa et al., 2021; Ktena et al., 2024; Movva et al., 2024): in particular, it can be used to estimate gaps in performance between groups, such as disparities in accuracy across race and gender  (Ji et al., 2020).
We conduct our analysis in the context of critical outcome prediction on MIMIC-IV, a task with documented prediction disparities (Movva et al., 2023). Like before, we still (1) acquire classifier scores for all patients using each classifier, and (2) fit the mixture model to classifier scores over the entire sample of labeled and unlabeled data. We then produce subgroup-specific performance estimates by using the empirical scores $\mathbf{s}$ within each subgroup; that is, we sample ground truth labels $y$ according to our estimated $p(y|\mathbf{s})$ for only those $\mathbf{s}$ observed among, for example, female patients. We then compare our estimated subgroup metric to the the ground truth metric evaluated on a large held-out sample for the given subgroup. We perform this analysis with respect to three categories of demographic groups: age, sex, and race/ethnicity. When there is no labeled data for a given subgroup, Labeled estimates subgroup-specific performance as global performance.
We report each method’s reduction in estimation error relative to labeled data (averaging over metrics and subgroups) for each demographic category in Table 2, comparing to all baselines which can estimate the four performance metrics we average over (accuracy, ECE, AUC, and AUPRC). SSME reduces metric estimation error by 5.3$\times$ on sex, 2.6$\times$ on race, and 2.4$\times$ on age relative to labeled data, and to a greater extent than all baselines. SSME also outperforms all baselines on all individual metrics except for AUC, for which Bayesian-Calibration reduces estimation error by 3.2$\times$ as compared to 2.2$\times$ for SSME. Bayesian-Calibration is particularly well-suited to estimating AUC because it assumes monotonicity when mapping $s$ to $y$; i.e., $P(y=1|s^{(i)})\geq P(y=1|s^{(j)})$ when $s^{(i)}\geq s^{(j)}$. When the classifiers in question have high AUCs — as the critical outcome classifiers do — this is a useful assumption to make.
6.4 Impact of classifier set characteristics on SSME’s performance
We now characterize SSME’s performance in a semi-synthetic setting, in which we can assess how characteristics like classifier accuracy and calibration affect the performance of the proposed approach. To do so, we create sets of three classifiers based on the widely-used Adult dataset (Becker & Kohavi, 1996), where the task is to predict whether a person’s income is above \$50K. To create differences between the three classifiers in a set, we train them on random fixed-size samples of 100 labeled examples from different portions of the dataset, partitioned based on age. In doing so, our semi-synthetic classifier sets mimic how training data for different real-world classifiers can differ in demographically meaningful ways. We repeat this procedure to produce 500 sets of three classifiers, where sets differ in the training data provided to each classifier. Our procedure naturally produces random variation in classifier properties, like accuracy and calibration, which we can use to study how well SSME performs. For additional experimental details, refer to Appendix B.3.
Classifier accuracy
We find that as the average accuracy among classifiers in the set increases, SSME’s performance estimation error decreases (Fig. 4, left). This trend holds across each of the four metrics we consider (see Figures S1 and S2 for additional plots) and can be attributed to how more accurate classifiers produce more separable components of the mixture model we aim to estimate. More accurate classifiers allow SSME to better estimate the ground truth label $y$ for unlabeled examples. Classifier accuracy directly impacts the separation of mixture model components, which existing work has shown to result in more accurate parameter estimation (Redner & Walker, 1984).
Classifier calibration
Similarly, when the classifier set grows more calibrated on average, SSME’s estimation error decreases. We can attribute this behavior to how poorly calibrated classifiers result in worse initialization for the mixture model, because we initialize component assignments for the unlabeled examples based on the mean prediction across classifiers.
Classifier quantity
Conditional on each of the classifier set characteristics discussed (accuracy and calibration), increasing the number of classifiers reduces performance estimation error. Our results suggest that increasing the number of classifiers can sometimes be more beneficial than acquiring a set of more accurate or better calibrated classifiers. Given the widespread availability of pretrained classifiers, this is a promising path towards more accurate evaluations.
Figure 4: Impact of classifier set characteristics ($\mathbf{n_{\ell}=20}$, $\mathbf{n_{u}=1000}$). We plot SSME’s performance as a function of average classifier set accuracy (left) and average classifier set calibration (right) among a set of semi-synthetic classifier sets, grouped into equal-width bins. More accurate and better calibrated classifier sets produce more accurate performance estimates, as expected. Adding classifiers to the set (dashed) can improve performance estimation to a greater extent than improving the average accuracy of the classifiers in the set: for example, adding two more classifiers typically produces a greater improvement than increasing average classifier accuracy by 5%.

### v3

@@ 4.1 Datasets and Classifier Sets
We select datasets and classifier sets to be realistic and diverse, capturing multiple modalities (EHRs, text, and graphs), domains (healthcare, content moderation, chemistry), and architectures (logistic regressions, graph neural networks, large language models). We report ground truth metrics for the binary and multiclass classifiers in Tables S1 and S2 respectively. We also include a detailed description of each dataset and classifier set in Appendix B.1. We summarize each dataset and differences between classifiers in the associated classifier set below.
We evaluate SSME on five classification datasets: (1) MIMIC-IV (johnson2020mimic, ), a dataset of Boston-area electronic health records (EHRs) with three hospitalization outcomes (critical outcome, ED revisit within 30 days, and hospital admission) with classifiers drawn from prior works (movva_coarse_2023, ); (2) CivilComments (borkan2019nuanced, ), a dataset of social media comments flagged as “toxic" or not by human annotators, with pretrained classifiers downloaded from the WILDS benchmark (koh2021wilds, ); (3) OGB-SARS-CoV (hu2020open, ), a molecular property prediction task, with pretrained classifiers downloaded from WILDS; (4) MultiNLI, a natural language task, with classifiers drawn from the SubpopBench dataset (yang2023change, ); and (5) AG News (zhang2015character, ), a news classification task, using open-source and closed-source LLMs as classifiers.
@@ 4.2 Baselines
We compare SSME to eight baselines that use both labeled and unlabeled data to estimate performance (with the exception of Labeled, the standard approach to classifier evaluation, which only uses examples for which labels are available). The remaining baselines cover a broad range of approaches one could adapt to semi-supervised evaluation and fall into three categories. For a full description of each baseline’s implementation, please refer to Appendix B.2.
The first category fits a model of $P(y,\mathbf{s})$ using labeled data alone, then applies it to the entire dataset . Each baseline does so in a different way: Pseudo-Labeled (PL) fits a logistic regression to predict the true label from classifier scores; Bayesian-Calibration (BC) (ji_can_2020, ) fits a Bayesian model to re-calibrate each classifier’s predicted probabilities; AutoEval (boyeau2024autoeval, ) learns a function to debias classifier predictions. The second category estimates the performance of individual classifiers. SPE welinder_lazy_2013 fits a parametric mixture model to a single classifier’s predictions. Active-Testing (kossen2021active, ) uses active learning to select examples to label out of a pool of unlabeled examples based on a single classifier’s predictions and evaluates the classifier on those examples. The last category of baselines is drawn from the multiple annotator literature, where annotations are assumed to be discrete. We compare to Dawid-Skene (dawid_maximum_1979, ), which accepts discrete annotations from each classifier to estimate the latent true label for each example. We also compare to Majority-Vote, which ensembles classifier predictions by performing an accuracy-weighted aggregation of classifier predictions.
We also provide a comparison to completely unsupervised approaches (including ensembling where the “ground truth" label is drawn based on the average classifier score) in Sec. D.10, and discuss connections to weak supervision in Sec. D.9.
@@ 4.3 Evaluation
We evaluate SSME’s ability to estimate four performance metrics for binary classifiers: accuracy, area under the receiver operating characteristic curve (AUC), area under the precision-recall curve (AUPRC), and the expected calibration error (ECE). For multi-class problems, we evaluate accuracy and top-label calibration error (gupta2022top, ).
We partition each dataset into three splits: the classifier training split (used to train the classifiers whose performance SSME estimates), the estimation split (used to fit SSME and estimate classifier performance), and the evaluation split (used to produce a held-out, ground-truth measure of classifier performance). All splits are sampled from the same distribution, except when estimating subgroup-specific performance, where the evaluation split contains a subset of the test distribution.
To evaluate metric estimates, we measure the absolute error of the estimated metric, computed using the estimation split, compared to the true metric, computed on the held-out evaluation split (averaging over classifiers in the set). The estimation split consists of either 20, 50, or 100 labeled examples and 1000 unlabeled examples across all experiments. The size of the evaluation split is on the order of thousands of labeled examples and varies by task (see Appendix B.1 for exact split sizes). To quantify uncertainty, we compute 95% confidence intervals using the standard error of the mean across 50 random data splits (±1.96 × SEM). In line with prior work, we report rescaled estimation error for each metric (where errors are relative to using labeled data alone), allowing us to standardize the scale of errors across datasets and metrics (garg_leveraging_2022, ).
@@ 5 Results
5.1 SSME produces more accurate performance estimates than prior work
We now compare SSME to eight baselines in terms of its ability to estimate classifier performance on five binary tasks. All figures report rescaled metric estimation error (RMAE; lower is better) and reflect performance estimation using 20 labeled examples and 1000 unlabeled examples. Results are consistent across additional values of $n_{\ell}$ (50 and 100; see results in D.2), although labeled data grows more competitive (as expected) with larger labeled dataset sizes.
Figure 2: Metric estimation error on binary tasks ($\mathbf{n_{\ell}=20}$, $\mathbf{n_{u}=1000}$). Each point plots the rescaled mean absolute error (RMAE) across runs, where 1.0 (dashed line) represents the RMAE when using labeled data alone and lower is better. SSME (gray) achieves lower estimation error than the baselines, averaging across metrics and 50 runs, reported for five binary tasks (y-axis). Results reporting mean absolute error (unscaled) and 95% CIs are in Tables S3-S6.
@@ 5.1 SSME produces more accurate performance estimates than prior work
[…]
Our results are also encouraging in absolute terms. With 20 labeled examples and 1000 unlabeled examples, SSME estimates accuracy within 1.5 percentage points (averaging across tasks). The closest baseline estimates accuracy within 3.4 percentage points. Results comparing SSME to the next best baseline on other metrics (1.9 vs 3.8 on ECE; 3.6 vs 4.3 on AUC; 8.5 vs 10.2 on AUPRC) confirm that SSME not only achieves more accurate classifier performance estimation compared to prior work, but that the resulting performance estimates are reasonably close to the ground truth.
Consistent with our theoretical results, SSME estimates performance more accurately because it makes use of all available information: multiple classifiers, continuous scores over all classes, and unlabeled data. Dawid-Skene and AutoEval discretize classifier scores (although AutoEval does make use of classifier confidence associated with each discrete prediction). Pseudo-Labeling, Bayesian-Calibration, and AutoEval each learn a mapping from s to $y$ using only the labeled data and apply that mapping to the unlabeled data (rather than learning from labeled and unlabeled data together). By jointly learning across both labeled and unlabeled data, SSME is able to generalize much better in cases where there aren’t enough labels to estimate the joint distribution of classifier scores and labels from labeled examples alone. Finally, Bayesian-Calibration and SPE learn from a single classifier’s scores, and do not learn from multiple classifiers at once.
@@
[…]
@@
[…]
Another way to quantify SSME’s benefit is to measure the amount of labeled data required to match SSME’s performance, or the effective sample size (ESS), as introduced by prior work (boyeau2024autoeval, ) (see Appendix B.3 for implementation details). With access to 20 labeled examples and 1000 unlabeled examples, SSME achieves an average ESS of 539 labeled examples for estimating ECE (averaging over tasks). In contrast, the next best approach achieves an ESS of 110 labeled examples.
@@
To validate the benefit of fitting the mixture model to multiple classifiers simultaneously, we compare to an ablated version of SSME fit on a single classifier at a time, SSME-M. SSME-M estimates the classifier-specific marginal distribution of $P(y|\textbf{s})$, and uses this estimate to evaluate the classifier in question. Doing so results in worse performance estimates across metrics, tasks, and amounts of labeled data (Tables S3- S6) relative to our full model. This agrees with our theoretical findings: each classifier provides distinct information about the ground truth label for a given example, which SSME is able to use.
Supplementary results
The supplement contains additional results. We report mean absolute error, including standard errors and additional $n_{\ell}$, for accuracy, ECE, AUC, and AUPRC in Tables S3- S6 (broken down by classifier in Tables S7- S10). We report the sensitivity of SSME to kernel choice in Table S13. Section D.11 provides a computational cost analysis of SSME relative to the baselines. Finally, we provide synthetic experiments to substantiate our theoretical findings under violations of our assumptions in Section D.1.
Figure 3: Metric estimation error for LLM classifiers ($\mathbf{n_{\ell}=20}$, $\mathbf{n_{u}=1000}$). SSME consistently reduces estimation error relative to labeled data alone and does so more reliably than any baseline capable of estimating performance in multiclass settings.
5.2 Case study: Evaluating classifiers derived from large language models
@@
BC 0.45 0.33 0.41
SSME (Ours) 0.42 0.19 0.39
@@
Table 1: Subgroup-specific performance estimation ($\mathbf{n_{l}=20}$, $\mathbf{n_{u}=1000}$). SSME achieves the lowest rescaled metric estimation error (RMAE), averaging across metrics and subgroups.
Large language models (LLMs) are increasingly used to create classifiers (kant2018practical, ; sun2023text, ; tornberg2024large, ), with demonstrated potential to serve as data annotators (ziems2023can, ). However, evaluation of LLM-based classifiers is limited by the availability of labeled data. We study SSME’s ability to estimate performance in this setting.
[…]
5.3 Case study: Evaluating subgroup-specific performance
SSME can be applied to measure performance within demographic subgroups, a task central to assessments of algorithmic fairness (chen2021ethical, ). We conduct our analysis in the context of critical outcome prediction on MIMIC-IV, a task with documented prediction disparities (movva_coarse_2023, ), and a set of commonly studied demographic groups (age, sex, race/ethnicity). We produce subgroup-specific performance estimates by sampling ground truth labels $y$ according to our estimated $p(y|\mathbf{s})$ for only those $\mathbf{s}$ observed among a given demographic group.
[…]
6 Related Work
Our work unifies ideas from two literatures: methods to 1) evaluate a single classifier and 2) evaluate the accuracy of multiple discrete annotators in the presence of labeled and unlabeled data.
Classifier evaluation is a long-standing problem in machine learning. Early work on the use of unlabeled data for classifier evaluation made progress via assumptions about the available unlabeled data (for example, assumptions about class priors balasubramanian2011unsupervised ; donmez_unsupervised_2010 ; lu_characterizing_2023 ; conditionally independent features steinhardt_unsupervised_2016 ; well-calibrated classifiers guillory_predicting_2021 ; models of covariate shift chen_estimating_2022 ; lee2023unsupervised ). SSME avoids such assumptions by using both labeled and unlabeled data to learn a flexible semi-supervised mixture model of classifier predictions. Interest in unsupervised evaluation has surged in the context of large language models, where models are often judged by other models (zheng2023judging, ; huang2024empirical, ). These approaches, while promising, are tailored to text generation and not applicable to broader classification tasks. A parallel line of work aims to make use of both labeled and unlabeled data to estimate classifier performance. Existing methods operate on individual classifiers and assume specific parametric forms for the prediction distributions (welinder_lazy_2013, ; chouldechova_unsupervised_2022, ; ji_active_2021, ; ji_can_2020, ; feofanov2023random, ). Some works also focus on model selection in semi-supervised settings, though they do not explicitly estimate performance (li2024towards, ). SSME differs from these works by naturally capitalizing on multiple classifiers, continuous classifier scores, and unlabeled data. As our results show, doing so leads to improved estimates of performance.
Evaluation of multiple discrete annotators was first introduced by dawid_maximum_1979 , who proposed a method to estimate ground truth in the presence of multiple potentially noisy discrete annotations. Many subsequent works inherit Dawid-Skene’s strong assumption of class-conditional independence of annotator errors (parisi_ranking_2014, ; platanios_estimating_2017, ), including popular approaches in weak supervision (ratner_snorkel_2017, ; bach_learning_2017, ; fu2020fast, ), where annotators are instead user-provided labeling functions. Such an assumption is plausible in certain contexts, but does not naturally translate to sets of candidate classifiers, whose predictions are likely to be correlated. However, methods from this line of work are designed to estimate the accuracy of binary annotations; they do not exploit the continuous probabilities available in multi-classifier evaluation. While some work has made progress towards accommodating continuous predicted probabilities (nazabal_human_2016, ; pirs_bayesian_2019, ), their focus is optimal aggregation, in contrast to our own, which is evaluation. Recent works use a continuous notion of classifier confidence in conjunction with discrete annotations from each annotator (goh2022crowdlab, ; boyeau2024autoeval, ), but do not use the distribution of classifier scores over all classes.